Projection Admin 2/5



To do this evening/ was completed: Hey everyone, the Weekly Update has been sent out and there's some super important info in there regarding the Super Bowl this weekend, or rather, the cable connection setup we'll be utilizing in theatres 1, 2 and 3, so please take a look when you can. Jack, you'll be working during the game this weekend, so let us know if you have any questions (Alec will likely work the Oscars). - Lillian 

We've received an update about the Theater 2 server's tendency to load scheduled playlists and leave them paused at the start: Chris from DMS thinks simplifying/improving the fire alarm set up will resolve this. The issue was caused by the GPI trigger cue which is necessary for the server to respond to the fire alarm, so we won't remove it. Chris should be able to come in soon to work on this. -AL

Presentation Issues:

Preshow Clips/Playlist: Built the A1 playlists for Heart Eyes, and the staff screening playlists, which have been scheduled on the LMS and blocked on the calendar. Email with staff screening info is out. 
Added/scheduled more shows on the LMS. Built and scheduled the playlists for Pieces, SOF, scheduled In the Mood for Love, Devil in a Blue Dress. Added Spoons this weekend as an "event" on the LMS. - Lillian 

Trailer Updates:

Special Events/ AV set up: Glory is all set.

Set up cable in Theater 3 for the Superbowl on Sunday night. I put some notes in the weekly update about this, but you should use the Digital 4 input on the CP750 for audio, NOT Non-Sync. Projector channel is BluRay HD 1.78. -AL

Ordering/Receiving (posters & trailers): 

Ordering/Receiving (content): Downloaded and ingested the preshow for Heart Eyes
Ingested the feature Slumber Party Massacre
Ingested Paddington 3
Chased down the new keys for Conclave (and Flow). We received the keys for Conclave this afternoon.
Lastly, I am currently downloading the DCP of Soundtrack to a Coup D'etait in 4/6 to the DCPIII drive. It's downloading via Cyberduck. Please copy the DCP over to the share drive once it's done. Thanks!  - Lillian 


Outgoing/Shipping (content):

Print/video Inspection: Took a look at The Brutalist in T4 just to make sure it's looking good. 
DCP of Slumber Party Massacre looks great. We'll tech all content with Desmond on Friday. - L

Asset tracking doc updated/checked?: I added the tracking for the SOF print in March, Into the Fire, and added all the rest of the titles/info for February and all of March. - Lillian 

Maintenance/cleaning: As per Will's note from last night about The Brutalist not ingesting to T4 due to storage, I spent a considerable amount of time today going through the stressed servers and carefully cleaning out the unnecessary bits of content - features/preshow/etc. A few of them will be a little pressed for a short period, but there's more wiggle room now. - Lillian 

Set 1.33 masking in Theater 1 and Theater 3 for upcoming Nickel Boys screenings. In order not to overwrite 1.66 in T3, I used the 2.20 preset and updated the cue in the playlist on the T3 server. -AL

More about the digital 4 input: it uses an optical cable, not RCA, and it sounds a lot louder than Non-Sync. This will be better for audio from HDMI connections in general. I ordered more optical cables and will make this a part of the HDMI-to-DVI setup in Theaters 1, 2, and 3. -AL

Cue sheets updated/checked (every shift):

Marquee updated?:

Posters updated?:

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Open 12/28

Projection Admin (and Close) 10/19

NPP Projection Report Friday 1/31 CLOSE Chris